OpenStudy (anonymous):

Find the equation of a line perpendicular to the given line: y = 3x – 1.

OpenStudy (anonymous):

so do u know what is the slope of y = 3x - 1?

OpenStudy (anonymous):

is it 3x over 1?

OpenStudy (anonymous):

but actually speaking you won't get a proper equation - bcoz if u still find the slope u do not have even one of the co-ordinates of the perpendicular line.

OpenStudy (anonymous):

well i dont what its asking me but here are the possible answers y = -1/3x – 1 y = 1/3x y = 1/3x – 1 y = -3x

OpenStudy (anonymous):

oh oh that way. Fine then. You know the slope of y = 3x-1 ?

OpenStudy (anonymous):

no it only says this Find the equation of a line perpendicular to the given line: y = 3x – 1.

OpenStudy (anonymous):

oh LOL that is FINE. but to find perpendicular equation, we need to find its slope. how would you do that?

OpenStudy (anonymous):

i know you add something to both sides

OpenStudy (anonymous):

well actually you DON'T. That is why i was asking you of the slope of y = 3x - 1.